# Cooperative distributed model predictive control for embedded systems: Experiments with hovercraft formations

### v1

@@ I Introduction
Model Predictive Control (MPC) is promising for robotics, because it explicitly accounts for actuator and safety constraints, interlaces motion planning with feedback control, and is applicable to output regulation, trajectory tracking, and path following [Katayama2023]. While early implementations of MPC were limited to slow dynamics because of the high computational demand, progress in real-time optimization schemes and hardware has enabled the application to fast and challenging systems such as aerial swarms and racing drones [Soria2021, Romero2022]. Moreover, MPC can be augmented with learning or data-driven approaches, expanding to applications where first-principles modeling is difficult [Hewing2020, Verheijen2023].
@@
Notation: Given a matrix $A$ and integers $i,j$, $[A]_{ij}$ is the component of $A$ at position $(i,j)$. Likewise, $[a]_{i}$ denotes the $i$th component of a vector $a$. The concatenation of vectors $a$ and $b$ into a column vector is denoted as $(a,b)$. Given an index set $\mathcal{S}$, the matrix $A=[A_{ij}]$ is the block matrix with entries $A_{ij}$ at block position $(i,j)$ for all $i,j\in\mathcal{S}$. Likewise, $B=\mathrm{diag}(R_{i})$ is the block diagonal matrix with entries $R_{i}$ at block position $(i,i)$ for all $i\in\mathcal{S}$. Given a vector $a$, $A=\mathrm{diag}(a)$ is the diagonal matrix where $[A]_{ii}=[a]_{i}$. The set $\mathbb{I}_{[0,N]}$ denotes the integers in range $[0,N]$ and $\mathbb{N}$ is the set of natural numbers.
@@ II Problem Statement
We consider the formation control of a robotic swarm $\mathcal{S}=\{1,\dots,S\}$ described by an undirected coupling graph $\mathcal{G}=(\mathcal{S},\mathcal{E})$. Neighboring robots are connected by edges $\mathcal{E}\subset\mathcal{S}\times\mathcal{S}$ in the graph and each robot can communicate bi-directionally with its neighbors $\mathcal{N}_{i}\doteq\{j\in\mathcal{S}|(i,j)\in\mathcal{E}\}$.
@@ II-A System Model and Formation Control
In a fixed global reference frame, $p_{i}=(p_{\mathrm{x},i},p_{\mathrm{y},i})$ denotes the position of the hovercraft center, $\varphi$ is the yaw angle, $v_{\mathrm{x},i}$ and $v_{\mathrm{y},i}$ are the center velocities, $\omega_{i}$ is the angular velocity, $u_{\mathrm{x},i}$ and $u_{\mathrm{y},i}$ are the center acceleration, and $u_{\varphi,i}$ is the angular acceleration.
@@ II-B Obstacle and Collision Avoidance
$\left\|(p_{\mathrm{x},i},p_{\mathrm{y},i})-(p_{\mathrm{x},j},p_{\mathrm{y},j})\right\|_{2}^{2}\geq{d^{2}_{\mathrm{min}}}$ (1a)
@@
with $d_{\mathrm{min}}>0$ for all $j\in\mathcal{N}_{i}$ and for all $i\in\mathcal{S}$. Here, $p_{\mathrm{x},\mathrm{obs}}(t)$ and $p_{\mathrm{y},\mathrm{obs}}(t)$ is the position of the obstacle at sampling instant $t$. This can be seen as a generalization of the on-demand collision avoidance and separating hyperplane strategies in [Luis2019, vanParys2017], because our dRTI scheme linearizes (1) in each control step, as described in the next section. To guarantee feasibility, we implement (1) as soft constraints [Luis2019].
@@ III Distributed Model Predictive Control Via Decentralized Real-Time Iterations
The decision variables ${\boldsymbol{x}}\doteq\left({x}[0],\dots,{x}[N]\right)$ and ${\boldsymbol{u}}\doteq({u}[0],\dots,{u}[N-1])$ with $u\doteq(u_{1},\dots,u_{S})\in\mathbb{R}^{n_{u}}$ are the predicted trajectories over the horizon $N$. Here and in the following, square brackets mark predicted variables and bold symbols denote trajectories. The objective penalizes the deviation from the setpoint $x_{\mathrm{d}}$ via the stage costs $\ell_{i}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u,i}}\rightarrow\mathbb{R}$,
@@
and the terminal penalties $V_{\mathrm{f},i}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}$,
@@
The initial condition on the input (2c) compensates for any delay up to the control sampling interval, because we solve OCP (2) for the input $u[1]$ to be applied in the next control step [Findeisen2006]. This is crucial, as solving OCP (2) with decentralized optimization and multiple communication rounds can be slow [Stomberg2023]. $\square$
To obtain a DMPC scheme with only neighbor-to-neighbor communication, we solve OCP (2) via the dRTI scheme from [Stomberg2024]. We introduce state copies ${w}_{ji}={x}_{j}$ for all $j\in\mathcal{N}_{i}$ which become additional decision variables of subsystem $i$ and replace ${x}_{j}$ in the coupled costs (2a) and state constraints (2f). Denote the predicted copy trajectories as ${\boldsymbol{w}}_{i}\doteq({\boldsymbol{w}}_{ji})_{j\in\mathcal{N}_{i}}$ with ${\boldsymbol{w}}_{ji}\doteq({w}_{ji}[0],\dots,{w}_{ji}[N])$. We collect the decision variables in the vector $z_{i}\doteq({\boldsymbol{x}}_{i},{\boldsymbol{u}}_{i},{\boldsymbol{w}}_{i})\in\mathbb{R}^{n_{i}}$ for all $i\in\mathcal{S}$ and rewrite OCP (2) as a partially separable Nonlinear Program (NLP)
@@
The functions $f_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}$, $g_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{g,i}}$, and $h_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{h,i}}$ are twice continuously differentiable and are composed of the components of OCP (2). That is, $f_{i}$ are the cost functions, $g_{i}$ include the initial conditions and system dynamics, and $h_{i}$ include the collision avoidance, obstacle avoidance, and box constraints. The coupling constraint (3c) with the sparse matrices $E_{i}\in\mathbb{R}^{n_{c}}\times\mathbb{R}^{n_{i}}$ enforces equivalence between the original and copied states.
@@
$\mathbb{E}\doteq\left\{\bar{z}\in\mathbb{R}^{n}\left|\;\sum_{i\in\mathcal{S}}E_{i}\bar{z}_{i}=c\right.\right\}.$
@@
1:Initialize: $t=0,u_{i}(0),{x}_{i,\mathrm{d}},z_{i}(0),\gamma_{i}(0)$, $k_{\mathrm{max}},l_{\mathrm{max}},\rho$
2:while true do
@@
5:     for $k=0,\dots,k_{\mathrm{max}}-1$ do
6:         Compute $\nabla f_{i}(z_{i}^{k}),\nabla g_{i}(z_{i}^{k}),\nabla h_{i}(z_{i}^{k})$
@@
8:         for $l=0,\dots,l_{\mathrm{max}}-1$ do
@@
15:              Set $\bar{\boldsymbol{w}}^{k,l+1}_{ji}=\bar{\boldsymbol{x}}^{k,l+1}_{j}$ for all $j\in\mathcal{N}_{i}$
@@ IV-A Hovercraft
Each hovercraft consists of racing drone hardware mounted on a 150150150\,150mm diameter foam disk, weighs 145145145\,145g, and is light enough to float on an air hockey table. Six propellers generate thrust in all directions within the x-y plane, allowing the system to be modeled as a point mass. Compared to the original system [Schwan2024], we here use an improved version that is easier to assemble and has enhanced computing capabilities. Instead of mounting the propellers on a 3D-printed ring, they are now directly attached to a printed circuit board, providing increased rigidity and improved power delivery. Moreover, the ESP32 microcontroller has been replaced by a Radxa Zero 3W, featuring a quad-core Arm Cortex A55 with clock speeds up to 1.61.61.6\,1.6GHz.
@@ IV-B State Estimation
An OptiTrack system measures the hovercraft position to millimeter accuracy using reflective markers mounted on the foam base. To estimate velocities and disturbances, we implement a continuous-time Extended Kalman Filter (EKF) with discrete measurements [OptimalStateEstimation2006]. A constant-disturbance model is employed, which is then integrated into the system dynamics of the OCP to achieve offset-free tracking. In particular, the EKF estimates the extended state $\xi_{i}\doteq(x_{i},d_{i})\in\mathbb{R}^{9}$, with the disturbance $d_{i}\doteq(d_{x,i},d_{y,i},d_{\varphi,i})\in\mathbb{R}^{3}$. The model (2b) is replaced by the disturbed system dynamics
$x_{i}[\tau+1]=A_{i}x_{i}[\tau]+B_{i}u_{i}[\tau]+B_{i}d_{i}\quad\forall\tau\in\mathbb{I}_{[0,N-1]},\$
@@ IV-C Communication
We rely on the Robot Operating System (ROS2) for communication between the onboard and offboard computing devices [Macenski2022]. As outlined in Figure 1, we have two experimental setups: 1) State estimation and DMPC algorithm run onboard the hovercraft, and only position measurements are transmitted externally. The hovercraft communicate via Wi-Fi using the reliable transport protocol provided by ROS2. We found that the best effort approach led to many messages being dropped. 2) In offboard experiments, state estimation and the DMPC algorithm run on two external Minisforum UM790 Pro computers connected via Ethernet. The nodes are distributed on the two computers such that communication in the DMPC algorithm has to go through the Ethernet layer to simulate proper delay, i.e., DMPC nodes of neighboring robots run on different machines.
@@ IV-D Distributed MPC Implementation and Design
The DMPC controller for each robot is implemented as a ROS node to facilitate communication. Vectors and matrices are stored with the Eigen library, and CasADi libraries are code-generated for the OCPs to evaluate derivatives [Andersson2019]. We use the sparse interface of the QP solver PIQP to solve the subsystem QPs in Line 9 of Algorithm 1 [Schwan2023]. To promote a timely delivery of control signals despite spikes in communication delay, each subsystem waits at most 252525\,25ms in each of Lines 11 and 14 of Algorithm 1 or until 757575\,75of the control sampling interval has elapsed. If not all expected messages have arrived by then, the respective step is completed asynchronously.
@@
We tuned $Q$ and $R$ in experiments to achieve sufficient damping of the linear-quadratic regulator-controlled system obtained as a byproduct from the Riccati equation. We tuned ADMM for fast convergence to the OCP minimizer in simulations and found that a range of penalty parameters works well. $\square$
@@
To ensure feasibility of the QPs passed to PIQP, we implement all state constraints as soft constraints. This is necessary because of inevitable disturbances in experiments as well as consensus errors due to the limited number of ADMM iterations. To this end, we add slack variables to all state inequality constraints and quadratic penalties on the slack variables to the objective. $\square$
@@ V Experiments
We test the DMPC scheme in experiments of varying difficulty including point-to-point transitions, trajectory tracking, and a dynamic obstacle. Video recordings of all experiments are available online. [^2] https://www.youtube.com/watch?v=fgF4SPxZ1VA To analyze strengths and weaknesses of the approach, we compare different OCP designs and, by switching between onboard and offboard computation, optimizer settings. To evaluate control performance, we consider the averaged closed-loop cost
@@ V-B Point-to-Point Transitions with Obstacle
Figure 3: Hovercraft motion in 2D plane for three maneuvers with onboard computation and sampling interval $\Delta t=50\,$ms. Solid and dotted lines show closed-loop and predicted trajectories, respectively. Circles are the current hovercraft positions and crosses mark the setpoints. The top row depicts an abrupt setpoint change with a reconfiguration from a line into a rectangular formation. The center row shows a switch in position between neighbors while avoiding collisions. In the bottom row, the four hovercraft cross the table from left to right while dodging a dynamic obstacle shown in red. The red dashed line is the future obstacle trajectory which is unknown to the hovercraft and which is plotted for better visualization of the experiment.

### v2

@@ I Introduction
Model Predictive Control (MPC) is promising for robotics, because it explicitly accounts for actuator and safety constraints, interlaces motion planning with feedback control, and is applicable to output regulation, trajectory tracking, and path following [Katayama2023]. While early implementations of MPC were limited to slow dynamics because of the high computational demand, progress in real-time optimization schemes and hardware has enabled the application to fast systems such as aerial swarms and racing drones [Soria2021, Romero2022].
@@
Notation: Given a matrix $A$ and integers $i,j$, $[A]_{ij}$ is the component of $A$ at position $(i,j)$. Likewise, $[a]_{i}$ denotes the $i$th component of a vector $a$. The concatenation of vectors $a$ and $b$ into a column vector is $(a,b)$. Given an index set $\mathcal{S}$, $(a_{j})_{j\in\mathcal{S}}$ is the vertical concatenation of the vectors $a_{j}$ for all $j\in\mathcal{S}$. The matrix $A=[A_{ij}]$ is the block matrix with entries $A_{ij}$ at block position $(i,j)$ for all $i,j\in\mathcal{S}$ and $B=\mathrm{diag}(R_{i})$ is the block diagonal matrix with entries $R_{i}$ at block position $(i,i)$ for all $i\in\mathcal{S}$. Given a vector $a$, $A=\mathrm{diag}(a)$ is the diagonal matrix where $[A]_{ii}=[a]_{i}$. The set $\mathbb{I}_{[0,N]}$ denotes the integers in range $[0,N]$ and $\mathbb{N}$ is the set of natural numbers.
@@ II Problem Statement
We consider the formation control of a robotic swarm $\mathcal{S}=\{1,\dots,S\}$ described by an undirected time-invariant coupling graph $\mathcal{G}=(\mathcal{S},\mathcal{E})$. Neighboring robots are connected by edges $\mathcal{E}\subseteq\mathcal{S}\times\mathcal{S}$ in the graph and each robot can communicate bi-directionally with its neighbors $\mathcal{N}_{i}\doteq\{j\in\mathcal{S}|(i,j)\in\mathcal{E}\}$.
@@ II-A System Model and Formation Control
In a fixed global reference frame, $p_{i}=(p_{\mathrm{x},i},p_{\mathrm{y},i})$ denotes the position of the hovercraft center, $\varphi$ is the yaw angle, $v_{\mathrm{x},i}$ and $v_{\mathrm{y},i}$ are the center velocities, $\omega_{i}$ is the angular velocity, $u_{\mathrm{x},i}$ and $u_{\mathrm{y},i}$ are the center acceleration, and $u_{\varphi,i}$ is the angular acceleration. Thus, given an initial state $x_{i,0}\in\mathbb{R}^{n_{x,i}}$, we model each robot as a time-invariant discrete-time double integrator
@@ II-B Obstacle and Collision Avoidance
$\left\|(p_{\mathrm{x},i},p_{\mathrm{y},i})-(p_{\mathrm{x},j},p_{\mathrm{y},j})\right\|_{2}^{2}$ $\geq{d^{2}_{\mathrm{min}}}\;\,\forall j\in\mathcal{N}_{i},$ (1a)
@@
with $d_{\mathrm{min}}>0$ for all $i\in\mathcal{S}$. Here, $p_{\mathrm{x},\mathrm{obs}}(t)$ and $p_{\mathrm{y},\mathrm{obs}}(t)$ is the obstacle position at sampling instant $t$.
@@ III Distributed Model Predictive Control Via Decentralized Real-Time Iterations
The decision variables ${\boldsymbol{x}}\doteq\left({x}[0],\dots,{x}[N]\right)$ and ${\boldsymbol{u}}\doteq({u}[0],\dots,{u}[N-1])$ with $u\doteq(u_{1},\dots,u_{S})\in\mathbb{R}^{n_{u}}$ are the predicted trajectories over the horizon $N$. Here and in the following, square brackets mark predicted variables and bold symbols denote trajectories. The objective penalizes the deviation from the setpoint $x_{\mathrm{d}}$ via the convex stage costs $\ell_{i}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u,i}}\rightarrow\mathbb{R}$,
@@
and the convex terminal penalties $V_{\mathrm{f},i}:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}$,
@@
The initial condition on the input (2c) compensates for any delay up to the control sampling interval, because we solve OCP (2) for the input $u[1]$ to be applied in the next control step [Findeisen2006]. This is crucial as solving OCP (2) with multiple communication rounds can be slow [Stomberg2023]. $\square$
To obtain a DMPC scheme with only neighbor-to-neighbor communication, we solve OCP (2) via the dRTI scheme from [Stomberg2024]. We introduce state copies ${w}_{ji}={x}_{j}$ for all $j\in\mathcal{N}_{i}$ which become additional decision variables of subsystem $i$ and replace ${x}_{j}$ in the coupled costs (2a) and state constraints (2f). Denote the predicted copy trajectories as ${\boldsymbol{w}}_{i}\doteq({\boldsymbol{w}}_{ji})_{j\in\mathcal{N}_{i}}$ with ${\boldsymbol{w}}_{ji}\doteq({w}_{ji}[0],\dots,{w}_{ji}[N])$. We collect the decision variables in the vector $z_{i}\doteq({\boldsymbol{x}}_{i},{\boldsymbol{u}}_{i},{\boldsymbol{w}}_{i})\in\mathbb{R}^{n_{i}}$ for all $i\in\mathcal{S}$ and rewrite OCP (2) as a non-convex partially separable Nonlinear Program (NLP)
@@
The functions $f_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}$, $g_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{g,i}}$, and $h_{i}:\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n_{h,i}}$ are three times continuously differentiable and are composed of the components of OCP (2). That is, $f_{i}$ are the cost functions, $g_{i}$ include the initial conditions and system dynamics, and $h_{i}$ include the collision avoidance, obstacle avoidance, and box constraints. The coupling constraint (3c) with the sparse matrices $E_{i}\in\mathbb{R}^{n_{c}}\times\mathbb{R}^{n_{i}}$ enforces equivalence between the original and copied states.
@@
$\mathbb{E}\doteq\left\{\bar{z}\in\mathbb{R}^{n}\left|\;\sum_{i\in\mathcal{S}}E_{i}\bar{z}_{i}=0\right.\right\}.$
@@
Remark 4 (Linearized collision avoidance)
Step 6 of Algorithm 1 linearizes the non-convex collision avoidance constraints (1), generalizing the on-demand collision avoidance and separating hyperplane strategies in [Luis2019, vanParys2017]. $\square$
@@
1:Initialize: $u_{i}(0),{x}_{i,\mathrm{d}},z_{i}(0),\gamma_{i}(0)$, $k_{\mathrm{max}},l_{\mathrm{max}},\rho$
2:for NMPC step $t=0,1,\dots$ do
@@
5:     for SQP iteration $k=0,\dots,k_{\mathrm{max}}-1$ do
6:         Compute $\nabla f_{i}(z_{i}^{k}),g_{i}(z_{i}^{k}),\nabla g_{i}(z_{i}^{k}),h_{i}(z_{i}^{k}),$    and $\nabla h_{i}(z_{i}^{k})$
@@
8:         for ADMM iteration $l=0,\dots,l_{\mathrm{max}}-1$ do
@@
15:               Set $\bar{\boldsymbol{w}}^{k,l+1}_{ji}=\bar{\boldsymbol{x}}^{k,l+1}_{j}\,\forall\,j\in\mathcal{N}_{i}$, $\bar{\boldsymbol{u}}^{k,l+1}_{i}={\boldsymbol{u}}^{k,l+1}_{i}$
@@ IV-A Hovercraft
Each hovercraft consists of racing drone hardware mounted on a 150150150\,150mm diameter foam disk, weighs 145145145\,145g, and can float on an air hockey table. Six propellers generate thrust in all directions within the x-y plane, allowing the system to be modeled as a point mass. Compared to the original system [Schwan2024], we here use an improved version that is easier to assemble and has enhanced computing capabilities. Instead of mounting the propellers on a 3D-printed ring, they are now directly attached to a printed circuit board, providing increased rigidity and improved power delivery. Moreover, the ESP32 has been replaced by a Radxa Zero 3W with a quad-core Arm Cortex A55 and clock speeds up to 1.61.61.6\,1.6GHz.
@@ IV-B State Estimation
An OptiTrack system measures the hovercraft position to millimeter accuracy using reflective markers mounted on the foam base. To estimate velocities and disturbances, we implement a continuous-time Extended Kalman Filter (EKF) with discrete measurements [OptimalStateEstimation2006]. A constant-disturbance model is employed, which is then integrated into the system dynamics of the OCP to achieve offset-free tracking. In particular, the EKF estimates the extended state $\xi_{i}\doteq(x_{i},d_{i})\in\mathbb{R}^{9}$, with the disturbance $d_{i}\doteq(d_{\mathrm{x},i},d_{\mathrm{y},i},d_{\varphi,i})\in\mathbb{R}^{3}$. The model (2b) is replaced by the disturbed system dynamics
$x_{i}[\tau+1]=A_{i}x_{i}[\tau]+B_{i}u_{i}[\tau]+B_{i}d_{i}\quad\forall\tau\in\mathbb{I}_{[0,N-1]},$
@@ IV-C Communication
We rely on the Robot Operating System (ROS2) for communication between the onboard and offboard computing devices [Macenski2022]. As outlined in Figure 1, we have two experimental setups: 1) The state estimation and DMPC algorithm run onboard the hovercraft, and only position measurements are transmitted externally. The hovercraft communicate via Wi-Fi using the reliable transport protocol provided by ROS2. We found that the best effort approach led to many messages being dropped. 2) In offboard experiments, the state estimation and DMPC algorithm run on two external Minisforum UM790 Pro computers connected via Ethernet. The nodes are distributed on the two computers such that communication in the DMPC algorithm has to go through the Ethernet layer to simulate proper delay, i.e., DMPC nodes of neighboring robots run on different machines.
@@ IV-D Distributed MPC Implementation and Design
A C++ implementation of Algorithm 1 is available online. [^2] https://github.com/PREDICT-EPFL/holohover The DMPC controller for each robot is implemented as a ROS node to facilitate communication. Vectors and matrices are stored with the Eigen library, and CasADi libraries are code-generated for the OCPs to evaluate derivatives [Andersson2019]. We use the sparse interface of the QP solver PIQP to solve the subsystem QPs in Line 9 of Algorithm 1 [Schwan2023]. To promote a timely delivery of control signals despite spikes in communication delay, each subsystem waits at most 252525\,25ms in each of Lines 11 and 14 of Algorithm 1 or until 757575\,75of the control sampling interval has elapsed. If not all expected messages have arrived by then, the respective step is completed asynchronously.
@@
We tuned $Q$ and $R$ in experiments for sufficient damping of the linear-quadratic regulator-controlled system obtained as a byproduct from the Riccati equation. We tuned ADMM in simulations and found a range of penalty parameters to work well. $\square$
@@
To ensure feasibility of the OCP, we implement all state constraints as soft constraints [Luis2019]. To this end, we add slack variables to all state inequality constraints and quadratic penalties on the slack variables to the objective. This is necessary because of inevitable disturbances in experiments as well as consensus errors due to the limited number of ADMM iterations. Consequently, violations of the state constraints may occur but are heavily penalized. $\square$
@@ V Experiments
We test the DMPC scheme in experiments of varying difficulty including point-to-point transitions, trajectory tracking, and a dynamic obstacle. Video recordings of all experiments are available online. [^3] https://www.youtube.com/watch?v=-ojLJUFMong To analyze strengths and weaknesses of the approach, we compare different OCP designs and, by switching between onboard and offboard computation, optimizer settings. To evaluate control performance, we consider the averaged closed-loop cost
@@ V-B Point-to-Point Transitions with Obstacle
Figure 3: Hovercraft motion in the 2D plane for three maneuvers with onboard computation and sampling interval $\Delta t=50\,$ms. Solid and dotted lines show closed-loop and predicted trajectories, respectively. Circles are the current hovercraft positions and crosses mark the setpoints. The top row depicts an abrupt setpoint change with a reconfiguration from a line into a rectangular formation. The center row shows a switch in position between neighbors while avoiding collisions. In the bottom row, the four hovercraft cross the table from left to right while dodging a dynamic obstacle shown in red. The red dashed line is the future obstacle trajectory which is unknown to the hovercraft and which is plotted for better visualization of the experiment.